Eight scores have an average of six. Scores of 15 and x increase the average to 7. Find x.
Answer by 428225(90)

8*6 = 48 (mean * amount numbers)
-------
Now add 15 and x to this number
-------
(48+15+x)/(10)=7
63+x=70
x=7